OpenStudy (anonymous):

A card is drawn from a well shuffled deck of 52 cards. Find P(drawing a face card or a 4). A face card is a king, queen, or jack. A. 4/13 B. 2/13 C. 12/13 D. 16

OpenStudy (triciaal):

this time probabilities this OR that

OpenStudy (anonymous):

nvm i fixed it

OpenStudy (anonymous):

this is probability, all of these are

OpenStudy (triciaal):

read options P(4) = 4/52 p (face)= 3*4/52 P (4 or face) = add so (4 + 12)/52 = 4/13

OpenStudy (triciaal):

yes A

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

what do u think dansih